OpenStudy (anonymous):

Bill and Steve decide to spend the afternoon at an amusement park enjoying their favorite activities, the water slide and the gigantic Ferris wheel. Their tickets are stamped each time they slide or ride. At the end of the afternoon they have the following tickets: Fun Time Amusements: Bill rode the water slide 3 times Bill rode rode the Ferris Wheel 3 times. His total: $17.70 Steve's Ticket- Steve rode the water slide two times. Steve rode the Ferris Wheel the three times. His total:$15.55

OpenStudy (anonymous):

What is the question?

OpenStudy (anonymous):

How much does it cost to the ride the Ferris Wheel? How much does it cost to ride the Water Slide?

OpenStudy (anonymous):

Using the information you have, you should come up with two equations using w, the cost of going down the Water Slide, and f, the cost of riding the ferris wheel. Can you figure out the two equations?

OpenStudy (anonymous):

Im kind of confused

OpenStudy (anonymous):

Okay, so you need two equations in order to solve this problem. You're trying to solve for two variables, or unknown amounts--- w (the cost of the water slide) and f (the cost of the ferris wheel ride). So, you have the information that, for example, Bill rode the ferris wheel three times and water slide three times, and the cost of all of his rides was 17.70. To turn that into equation form, 3w + 3f = 17.70 Does that make sense? Can you get an equation from Steve's information?

OpenStudy (anonymous):

Steve's equation: 2w + 3f = 155

OpenStudy (anonymous):

awesome! Now, you have to use "elimination" to figure out how much w and f each are. So... 3w + 3f = 17.70 - 2w + 3f = 15.55 ----------------------- ?

OpenStudy (anonymous):

Then I got , w+6f=33.25

OpenStudy (anonymous):

Subtract instead of adding! You're trying to eliminate f and only have w :)

OpenStudy (anonymous):

so... w=33.25 ?

OpenStudy (anonymous):

You've gotta subtract 15.55 from 17.70 too. 3w + 3f = 17.70 - 2w + 3f = 15.55 ----------------------- w + 0f = 2.15 w = 2.15

OpenStudy (anonymous):

don't stress, you're getting there! does that make sense, so far? The next step is putting the solution for w back into one of the equations to solve for f. 2w + 3f = 15.55 2(2.15) + 3f = 15.55 Can you get f by itself on one side of the equation?

OpenStudy (anonymous):

yes. I got my answer and f ended up being f=3.75

OpenStudy (anonymous):

yes!! nice job

OpenStudy (anonymous):

one more question :P The equations 5x+2y=48 and 3x+2y=32 represent the money collected from school concert ticket sales during two class periods. If x represents the cost for each adult ticket and y represents the cost for each student ticket, what is the cost for each adult ticket.?

OpenStudy (anonymous):

This is the exact same type of problem. You're trying to isolate x and completely get rid of y, so you need to use elimination. Does that make sense?

OpenStudy (anonymous):

yes , I suppose.

OpenStudy (anonymous):

Set it up in the same way we set up the other problem, in order to get rid of y, and just have the variable x, which we want to figure out the answer to. Remember it's subtraction! 5x+2y=48 - 3x+2y=32 ----------------